# Anisotropy of emergent large-scale dynamics in forced stratified shear flows

### v1

@@ 4 Discussion and perspective
[Acknowledgements] P.P.V. (i) thanks Roshan J. Samuel (TU Ilmenau, Germany) for early help with NekRS and (ii) gratefully acknowledges the support of Pembroke College, Cambridge, through a postdoctoral research associateship.
@@
[Author ORCIDs] P. P. Vieweg, https://orcid.org/$0$-0001-7628-9902
C. P. Caulfield, https://orcid.org/$0$-0002-3170-9480
@@
Appendix A Kinetic energy and buoyancy variance equations
All of our simulations are based on equations (3) and (4) as the evolution equations for the momentum and buoyancy, respectively. However, some quantities used in the main text are actually related to the evolution equations of kinetic energy $\bm{u}^{2}/2$ or buoyancy variance $b^{2}/2$ instead. These two (non-dimensional) equations are
$\frac{\partial}{\partial t}\frac{\bm{u}^{2}}{2}$ $=-\nabla\cdot\bm{J}_{u}-\varepsilon_{u}+F_{u}+B,$ (26)
$\frac{\partial}{\partial t}\frac{b^{2}}{2}$ $=-\nabla\cdot\bm{J}_{b}-\varepsilon_{b}+F_{b}$ (27)
and follow from taking the dot product of equation (3) with $\bm{u}$ or multiplying (4) by $b$. Both equations have a similar structure.
The dissipation rates $\varepsilon_{\Phi}$ have already been introduced in the main text in equations (17) and (18), whereas the vertical buoyancy advection or buoyancy flux $B$ has been introduced in equation (16). In addition, the $j$-th components of the flux terms $\bm{J}_{\Phi}$ are given by
$J_{u,j}$ $:=u_{j}\frac{u_{i}^{2}}{2}+u_{j}p-\frac{1}{\mathrm{Re}_{0}}\left[\frac{\partial}{\partial x_{j}}\frac{u_{i}^{2}}{2}+\frac{\partial}{\partial x_{i}}\left(u_{i}u_{j}\right)\right],$ (28)
$J_{b,j}$ $:=u_{j}\frac{b^{2}}{2}-\frac{1}{\mathrm{Re}_{0}\mathrm{Pr}}\frac{\partial}{\partial x_{j}}\frac{b^{2}}{2}$ (29)
whereas our additional source terms $F_{\Phi}$ are given by
$F_{u}:=u_{x}f_{u}\qquad\textrm{and}\qquad F_{b}:=bf_{b}.$ (30)
Note that we apply the Einstein summation convention in equations (28) and (29) for improved clarity. The volumetric forcings $f_{\Phi}$ have been defined in the main text in equations (6) and (7).
Equation (27) represents the evolution equation for the unscaled buoyancy variance. The evolution equation for the scaled buoyancy variance follows simply from a subsequent multiplication of this equation by $\mathrm{Ri}_{0}^{2}/\langle N^{2}\rangle_{A,t}$, see also again equations (18), (20), and (21).
Appendix B Resolving shear flows efficiently using spectral element methods
@@
Figure 9: Resolving shear flows. (a) Shear flows exhibit highly non-uniform profiles of dissipation $\varepsilon_{u}$ and, thus, the smallest dynamical scales $\eta_{\textrm{K}}$. (b) Spectral element methods allow for the adjustment of the height of spectral elements $h_{\mathrm{se}}$. The local resolution $dz$ follows from a subsequent spectral expansion of polynomial order $N$ within each element. Hence, adjusting $h_{\mathrm{se}}$ enables (c) the very efficient resolution of shear flows. Here, just like in figure 2, $L_{\mathrm{h}}=128$.
@@
Appendix C Convergence of horizontal energy spectra
@@
In section 3.4, we have introduced Fourier energy spectra and explained that the location of their peak serves as a measure to quantify a length scale which is characteristic of the large-scale dynamics. That analysis has been conducted for our largest square domain at $L_{\mathrm{h}}=512$ and underlined by even larger but non-square domains, see particularly figure 7.

### v2

@@ 5 Conclusions
[Acknowledgements] P.P.V. (i) thanks Roshan J. Samuel (TU Ilmenau, Germany) for early help with NekRS and (ii) gratefully acknowledges the support of Pembroke College, Cambridge, through a postdoctoral research associateship. The constructive and insightful comments of three anonymous referees have led to significant improvements in this document, and their helpful and thoughtful input is gratefully acknowledged.
@@
[Author ORCIDs] P. P. Vieweg, https://orcid.org/0000-0001-7628-9902
C. P. Caulfield, https://orcid.org/0000-0002-3170-9480
@@
Appendix A Resolving shear flows efficiently using spectral element methods
@@
Figure 10: Resolving shear flows. (a) Shear flows exhibit highly non-uniform profiles of dissipation $\varepsilon_{u}$ and, thus, the smallest dynamical scales $\eta_{\textrm{K}}$. (b) Spectral element methods allow for the adjustment of the height of spectral elements $h_{\mathrm{se}}$. The local resolution dzdz follows from a subsequent spectral expansion of polynomial order NN within each element. Hence, adjusting $h_{\mathrm{se}}$ enables (c) the very efficient resolution of shear flows. Here, just like in figure 2, $L_{\mathrm{h}}=128$.
@@
Appendix B Convergence of horizontal energy spectra
@@
In section 3.4, we have introduced Fourier energy spectra and explained that the location of their peak serves as a measure to quantify a length scale which is characteristic of the large-scale dynamics. That analysis has been conducted for our largest square domain at $L_{\mathrm{h}}=\Gamma_{\textrm{h, max}}=512$ and underlined by even larger but non-square domains, see particularly figure 7.